# High-order WENO finite-difference methods for hyperbolic nonconservative systems of Partial Differential Equations

### v1

@@ 3 Problems with source terms and well-balanced property
where $H(x)$ is a known function. The well-balanced property of the methods is related to the preservation of the stationary solutions $U^{*}$ of the system, which satisfy the equations
$A(U^{*})U^{*}_{x}=S(U^{*})H_{x}.$
Two strategies are considered here to discretize the source term:
3.1 Strategy 1
The first strategy consists in writing (24) in the form (5) as follows:
$W_{t}+\widetilde{A}(W)W_{x}=0,$
@@ 3.1 Strategy 1
and then the general strategy is applied to derive high-order WENO finite-difference methods. To do this, a family of paths
@@
has to be chosen. If the artificial unknown $H$ is dropped, the numerical method writes again as (14)-(15)-(16) where now
$D_{j,k}=\int_{0}^{1}A(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{U}(s;W_{j},W_{k})\,ds-\int_{0}^{1}S(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{H}(s;W_{j},W_{k})\,ds,$ (25)
where the notation $W_{j}=[U_{j},H(x_{j})]^{T}$ has been used. Moreover, if a Roe linearization is available, i.e. if a matrix-valued function $A_{\widetilde{\Psi}}(W_{L},W_{R})$ and a vector-valued function $S_{\widetilde{\Psi}}(W_{L},W_{R})$ exist such that
  • $A_{\widetilde{\Psi}}(W_{L},W_{R})$ has real different eigenvalues;
  • $A_{\widetilde{\Psi}}(W,W)=A(U)$ for all $W=[U,H]^{T}$;
@@
    $A_{\widetilde{\Psi}}(W_{L},W_{R})(U_{R}-U_{L})-S_{\widetilde{\Psi}}(W_{L},W_{R})(H_{R}-H_{L})$
    $=\int_{0}^{1}A(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{U}(s;W_{j},W_{k})\,ds-\int_{0}^{1}S(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{H}(s;W_{j},W_{k})\,ds,$
then the numerical method can be written as (14)-(15)-(16) with
$D_{j,k}=A_{j,k}(U_{k}-U_{j})-S_{j,k}(H(x_{k})-H(x_{j})),$ (26)
@@
$A_{j,k}=A_{\widetilde{\Psi}}(W_{j},W_{k}),\quad S_{j,k}=S_{\widetilde{\Psi}}(W_{j},W_{k}).$
Accordingly, if the LF splitting has been chosen, one has then
$D^{\pm}_{j,k}=A^{\pm}_{j,k}(U_{k}-U_{j})-S^{\pm}_{j,k}(H(x_{k})-H(x_{j})),$ (27)
where
$A^{\pm}_{j,k}=\frac{1}{2}\left(A_{j,k}\pm\alpha I\right),\quad S^{\pm}_{j,k}=\frac{1}{2}S_{j,k}.$
If the upwind approach is followed, the fluctuations are computed as follows: given two indices $j,k$, first the coordinates $\{\alpha_{j,k;l}\}_{l=1}^{N}$ of $U_{k}-U_{j}-A_{j,k}^{-1}S_{j,k}(H(x_{k})-H(x_{j}))$ in the basis of eigenvectors of the Roe matrix $A_{j,k}$ are computed:
@@
The well-balanced property of methods desiged using this first strategy is given by the following result:
Proposition 2.
Let $U^{*}(x)$ be a stationary solution of system (24). Then, if the matrix splitting is such that
$D_{i,j}=0\implies D^{\pm}_{i,j}=0,\quad\text{for }j=i-k,\dots,i+k,$ (29)
and the chosen family of paths $\widetilde{\Psi}$ is such that, given $x_{L},x_{R}\in\mathbb{R}$ the functions
$s\in[0,1]\to\Psi_{U}\left(s;\left[\begin{array}[]{c}U^{*}(x_{L})\\ H(x_{L})\end{array}\right];\left[\begin{array}[]{c}U^{*}(x_{R})\\ H(x_{R})\end{array}\right]\right)$
@@
$x\in[x_{L},x_{R}]\to U^{*}(x)$
define the same curve in $\Omega$, then the numerical method (14)-(15)-(16)-(25) is well-balanced for $U^{*}$, i.e. $\{U^{*}(x_{i})\}$ is an equilibrium of the ODE system (14).
Proof.
Let us check that $\{U^{*}(x_{i})\}$ is an equilibrium for the ODE system defined by the semi-discretized numerical method (14), i.e. that
$\widehat{D}_{i-1/2}^{+}+\widehat{D}_{i+1/2}^{-}=0,$
for $W_{i}=[U^{*}(x_{i}),H(x_{i})]^{T}$. In effect, we have
$D_{i,j}$ $=$ $\int_{0}^{1}A(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{U}(s;W_{j},W_{k})\,ds-\int_{0}^{1}S(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{H}(s;W_{j},W_{k})\,ds$
@@
$=$ $0,$
where a change of parameter has been applied to obtain the second equality and the fact that $U^{*}$ is a stationary solution has been used in the third one. Then, due to (29), from (15)-(16) we derive
$\widehat{D}_{i-1/2}^{+}=\widehat{D}_{i+1/2}^{-}=0,$
what ends the proof. ∎
Observe that (29) is satisfied for the upwind approach but it is not for the LF splitting is used: in effect, if $D_{i,j}=0$ one has
$D^{\pm}_{i,j}=\pm\frac{\alpha}{2}(U_{j}-U_{i})$
which is in general different from 0. Therefore, Strategy 1 for well-balancing cannot be combined with the choice of LF splitting.
3.2 Strategy 2
Strategy 2 extends to the nonconservative system the technique proposed in [38] for systems of balance laws: to obtain a fully well-balanced method, let us denote by $U^{*}_{i}$ the stationary solution satisfying
$U^{*}_{i}(x_{i})=U_{i},$
if it exists. Then, the equalities
$S(U_{i})H_{x}(x_{i})=S(U^{*}_{i}(x_{i}))H_{x}(x_{i})=A(U^{*}(x_{i}))(U^{*}_{i})_{x}(x_{i})$
can be used to approximate the source term, which leads to the approximate equation
$U_{t}+A(U)U_{x}-A(U^{*}_{i})(U^{*}_{i})_{x}(x_{i})=0,$
that is exactly at $x_{i}$. The idea is then to consider the numerical method (14) where now
@@ 3.2 Strategy 2
$+\mathcal{R}^{R}(D^{-}_{i-k,i}-D^{*,-}_{i-k,i},\dots,D^{-}_{i+k,i}-D^{*,-}_{i+k,i}).$
with
$D^{*,\pm}_{j,k}=\int_{0}^{1}A^{\pm}(\Psi(s;U_{i}^{*}(x_{j}),U_{i}^{*}(x_{k}))\partial_{s}\Psi(s;U^{*}_{i}(x_{j}),U_{i}^{*}(x_{k}))\,ds,$ (32)
or
@@
if a Roe linearization of $A$ is available.
@@
The numerical method (14)-(15)-(16) is fully well-balanced, i.e. $\{U^{*}(x_{i})\}$ is an equilibrium of the ODE system (14) for every stationary solution $U^{*}$.
Proof.
It is clear that
$U^{*}_{i}\equiv U^{*}$
is the stationary solution that satisfies
$U^{*}_{i}(x_{i})=U^{*}(x_{i})$
for all $i$. Therefore,
$D^{*,\pm}_{j,k}=D^{\pm}_{j,k}$
so that
$\widehat{D}^{+}_{i+1/2}=\widehat{D}^{+}_{i-1/2}=0,$
what ends the proof. ∎
Remark.
Observe that, to prove the result, it has been assumed that the matrix splitting used to approximate $A(U^{*}_{i})\partial_{x}U^{*}_{i}$ is exactly the same as the one used to approximate $A(U)U_{x}$. In particular, the same constant $\alpha$ has to be used in the LF splitting.
This strategy can be easily adapted to design numerical methods that only preserve a given $m$-parameter family of stationary solutions
$U^{*}(x;C_{1},\dots,C_{m})$
with $m<N$. The idea is to select the function of this family
$U^{*}_{i}(x)=U^{*}(x;C_{i,1},\dots,C_{i,m})$
whose parameters $C_{i,j}$, $j=1,\dots,m$ are such that
$u^{*}_{i;j_{l}}(x_{i})=u_{i;j_{l}},\quad l=1,\dots,m,$
where $u^{*}_{i;j}$ and $u_{i;j}$ represent respectively the $j$th component of $U^{*}_{i}(x)$ and $U_{i}$; and $j_{1},\dots,j_{m}$ are a set of pre-determined indices whose selection depends on the particular problems. The numerical method is in this case as follows:
$U_{i}^{\prime}+\frac{1}{\Delta x}\left(\widehat{D}^{-}_{i+1/2}+\widehat{D}^{+}_{i-1/2}\right)=(S(U_{i})-S(U^{*}_{i}(x_{i})))H_{x}(x_{i}),$ (34)
together with (30)-(31)-(32). It can be trivially checked that it preserves the stationary solutions of the $m$-parameter family.

### v2

@@ 2.3 Accuracy of the methods
The symmetry condition (38) is satisfied by the family of straight segments
$\Psi(s;U,V)=U+s(V-U).$
In effect
[…]
$=$ $-\int_{0}^{1}A(\Psi(s;U,V))\partial_{s}\Psi(s;U,V)\,ds.$
Therefore, in the particular case of system (4), the definition (24), based on the choice of straight segments, leads to a high-order method. On the other hand, (38) is not satisfied for (7) and, it will be seen in Section 6.1 that the method corresponding to (25) is only first-order accurate, what shows that this condition is necessary as well.
According to the proof of Proposition 1, the numerical method can be interpreted as follows: the PDE system is first formally rewritten as the system of balance laws
$\partial_{t}{U_{i}}+\partial_{x}\mathcal{F}^{L}_{i}=0,$ (41)
with
$\mathcal{F}^{L}_{i}(x,t)=G_{i}^{t}(x),$
where $G_{i}$ is the function given by (40); then WENO reconstructions are applied to the generalized flux function $\mathcal{F}^{L}_{i}$. In the particular case of a system of the form (11) it can be easily checked that this is equivalent to reconstructing the generalized flux function
$\mathcal{F}^{L}=F+\mathcal{B}^{L}_{i},$
where
$\mathcal{B}^{L}_{i}(x,t)=\int_{0}^{1}B(\Psi(s;U(x_{i},t),U(x,t))\partial_{s}\Psi(s;U(x_{i},t),U(x,t))\,ds,$
which is defined for every $i$, while in the global-flux approach the generalized flux function to be reconstructed is
$\mathcal{F}^{G}=F+\mathcal{B}^{G},$
where
[…]
is globally defined. This is why it was said above that a local flux approach is followed here. Following this approach, $\mathcal{B}^{L}_{i}$ is approximated at the node points of the stencil as follows
$\mathcal{B}^{L}_{i}(x_{j},t)\approx B_{\Psi}(U_{i},U_{j})(U_{j}-U_{i}),\quad j\in\mathcal{S}_{i},$
where $B_{\Psi}$ is the linearization of $B$ used in the Roe matrix, While in the case of the global flux approach (similar to the approach in the finite volume method as [10]), $\mathcal{B}^{G}$ is numerically approximated at the nodes using a recursive formula such that
$\mathcal{B}^{G}_{0}=0;\quad\mathcal{B}^{G}_{i+1}=\mathcal{B}^{G}_{i}+\Delta x\sum_{l=0}^{M}\alpha_{l}B(U_{i}^{l})D_{x}U_{i}^{l},\quad i=0,\dots,NP-1,$
where $\alpha^{l}$, $l=0,\dots,M$ are the weights of the selected quadrature form and $U_{i}^{l}$, $D_{x}U_{i}^{l}$, $l=0,\dots,M$ are high-order approximations of $U$ and $U_{x}$ at the quadrature points, $NP$ is the total number of discrete points. Therefore $M+1$ additional reconstructions (of state in this case) are necessary. Summing up, while in the local flux approach two flux WENO reconstructions per point are needed (to compute $\widehat{D}_{i-1/2}^{+}$ and $\widehat{D}_{i+1/2}^{-}$), in the global flux approach one flux WENO reconstruction per intercell and $M+1$ state reconstructions per cell are necessary. It means that the local flux approach requires $2NP$ flux reconstructions per stage of the ODE solver used for the temporal discretization while the global flux approach requires $(M+2)NP$. On the other hand, if the order of the WENO reconstruction is $12k+$ 1, then $M$ has to be greater or equal than $k$ so that the numerical method preserves the order of the WENO reconstructions. Therefore, the number of reconstructions in the global flux approach is greater than $(k+2)NP$ compared to the $2NP$ reconstructions in the local flux approach.
@@ 3 Problems with source terms and well-balanced property
3.1 Well-balanced property
Let us first consider a system of the form (10) in which $\lambda=0$ is an eigenvalue of $A(U)$ for every $U\in\Omega$. The well-balanced property of the methods is related to the preservation of the stationary solutions $U^{*}$ of the system, which satisfy the equation
$A(U^{*})U^{*}_{x}=0.$
Observe that, $U^{*}_{x}$ is an eigenvector associated with the null eigenvalue for all $x$ such that $U^{*}(x)_{x}\not=0$. As an example, it can be easily checked that the stationary solutions of (4) are the set of functions
$U^{*}(x)=\left[\begin{array}[]{c}u^{*}(x)\\ v^{*}(x)\end{array}\right]\text{ s.t. }u^{*}(x)+v^{*}(x)=constant.$ (42)
The method described in Section 2.2 has then the well-balanced property given by the following results the proof of which is trivial:
Proposition 2.
Let $U^{*}(x)$ be a stationary solution of system (10). If, for every $x_{L}<x_{R}$ one has
$A_{\Psi}(U^{*}(x_{L}),U^{*}(x_{R}))(U^{*}(x_{R})-U^{*}(x_{L}))=0$ (43)
and the selected matrix-splitting is such that
$A_{\Psi}(U,V)(V-U)=0\implies A^{\pm}_{\Psi}(U,V)(V-U)=0,$ (44)
then the numerical method (19)-(20)-(21)-(28) is well-balanced for $U^{*}$, i.e. $\{U^{*}(x_{i})\}$ is an equilibrium of the ODE system (19).
Observe that (44) is satisfied for the Upwind splitting approach, as can be easily deduced from (32), but not for the LF splitting: in effect, in this case one has
$A_{\Psi}(U,V)(V-U)=0\implies A^{\pm}_{\Psi}(U,V)(V-U)=\pm\frac{\alpha}{2}(V-U).$
Nevertheless, the modification of the identity matrix technique introduced in [21] can be applied to modify the splitting so that (43) is satisfied.
As an application of Proposition 2, it can be easily checked that the property (43) is satisfied for the Roe matrix $A_{\Psi_{1}}$ defined in (27) for every stationary solution (42) of System (4). Therefore, the choices of the family of straight segments and the Upwind splitting lead to a numerical method for (4) that is fully well-balanced (and high-order accurate). On the other hand, (43) is not satisfied for the Roe matrix $A_{\Psi_{2}}$.
Property (43) is discussed in [43] in relation with the well-balanced property of Roe methods. Let us only recall that, given a stationary solution $U^{*}$, this property is satisfied if the family of paths is such that, for all $x_{L},x_{R}\in\mathbb{R}$ with $x_{L}<x_{R}$, the functions
$s\in[0,1]\to\Psi\left(s;U^{*}(x_{L}),U^{*}(x_{R})\right)\in\Omega$
and
$x\in[x_{L},x_{R}]\to U^{*}(x)\in\Omega$
define the same curve in $\Omega$. In effect, if this is the case one has:
$A_{\Psi}(U,V)(V-U)$ $=$ $\int_{0}^{1}A(\Psi(s;U,V))\partial_{s}\Psi(s;U,V)\,ds$
[…]
$=$ $0,$
where a change of parameter has been applied to obtain the second equality and the fact that $U^{*}$ is a stationary solution has been used in the third one. In particular, if the family of straight segments is chosen, the property (43) is satisfied for all stationary solutions such that the curve defined by $x\to U^{*}(x)$ lies in a straight line: this is the case of (4) whose stationary solutions (42) lie in a straight line of equation $u+v=constant$ in the $u,v$ plane.
3.2 Source terms
@@
where $H(x)$ is a known function, whose stationary solutions satisfy
$A(U^{*})U^{*}_{x}=S(U^{*})H_{x}.$ (46)
3.2.1 Strategy 1
The first strategy consists in writing (45) in the form (10) as follows (see [13, 43]):
$W_{t}+\mathcal{A}(W)W_{x}=0,$ (47)
@@ 3.2.1 Strategy 1
and then the strategy described in Section 2.2 is applied to (47). To do this, a family of paths
@@
satisfying (38) (as, for instance, the family of straight segments) and a Roe linearization have to be chosen first. As in [43], let us assume that a Roe matrix of the form
$\mathcal{A}_{\widetilde{\Psi}}(W_{L},W_{R})=\left[\begin{array}[]{c|c}A_{\widetilde{\Psi}}(W_{L},W_{R})&-S_{\widetilde{\Psi}}(W_{L},W_{R})\\ \hline\cr 0&0\end{array}\right]$
is available, where
  • $A_{\widetilde{\Psi}}(W_{L},W_{R})$ has $N$ real different eigenvalues $\lambda_{i}(W_{L},W_{R})$, $i=1,\dots,N$;
  • $A_{\widetilde{\Psi}}(W,W)=A(W)$ for all $W=[U,H]^{T}$;
  • $S_{\widetilde{\Psi}}(W,W)=S(W)$ for all $W=[U,H]^{T}$;
@@
    $A_{\widetilde{\Psi}}(W_{L},W_{R})(U_{R}-U_{L})=\int_{0}^{1}A(\Psi_{U}(s;W_{L},W_{R}))\partial_{s}\Psi_{U}(s;W_{L},W_{R})\,ds;$
    $S_{\widetilde{\Psi}}(W_{L},W_{R})(H_{R}-H_{L})=\int_{0}^{1}S(\Psi_{U}(s;W_{L},W_{R}))\partial_{s}\Psi_{H}(s;W_{L},W_{r})\,ds.$
In this paragraph, the Upwind splitting is considered. Some algebraic computations show that the corresponding splitting is given by the matrices
$\mathcal{A}^{\pm}_{\widetilde{\Psi}}(W_{L},W_{R})=\left[\begin{array}[]{c|c}P^{\pm}_{\widetilde{\Psi}}(W_{L},W_{R})A_{\widetilde{\Psi}}(W_{L},W_{R})&-P^{\pm}_{\widetilde{\Psi}}(W_{L},W_{R})S_{\widetilde{\Psi}}(W_{L},W_{R})\\ \hline\cr 0&0\end{array}\right]$
where $P^{\pm}_{\widetilde{\Psi}}$ are the projection matrices defined as in (33).
If the trivial equation for the artificial unknown $H$ is removed, the numerical method can be written again as (19)-(20)-(21) where now
$D^{\pm}_{j,k}=P_{j,k}^{\pm}\left(A_{j,k}(U_{k}-U_{j})-S_{j,k}(H(x_{k})-H(x_{j}))\right),$ (48)
@@
$P_{j,k}^{\pm}=P^{\pm}_{\widetilde{\Psi}}(W_{j},W_{k}),\quad A_{j,k}=A_{\widetilde{\Psi}}(W_{j},W_{k}),\quad S_{j,k}=S_{\widetilde{\Psi}}(W_{j},W_{k}).$
Accordingly, the fluctuations can be computed as follows: given two indices $j,k$, first the coordinates $\{\alpha_{j,k;l}\}_{l=1}^{N}$ of $U_{k}-U_{j}-A_{j,k}^{-1}S_{j,k}(H(x_{k})-H(x_{j}))$ in the basis of eigenvectors of the Roe matrix $A_{j,k}$ are computed:
@@
Proposition 2 can be then applied to this particular case to show that, given $H(x)$, the numerical method is well-balanced for a stationary solution $U^{*}$ provided that (43) holds, i.e. if
$A_{\widetilde{\Psi}}(U^{*}(x_{L}),U^{*}(x_{R}))(U^{*}(x_{R})-U^{*}(x_{L}))=S_{\widetilde{\Psi}}(U^{*}(x_{L}),U^{*}(x_{R}))(H(x_{R})-H(x_{L}))$ (50)
for all $x_{L}<x_{R}$. This is the case if
$s\in[0,1]\to\widetilde{\Psi}\left(s;\left[\begin{array}[]{c}U^{*}(x_{L})\\ H(x_{L})\end{array}\right];\left[\begin{array}[]{c}U^{*}(x_{R})\\ H(x_{R})\end{array}\right]\right)$ (51)
@@
$x\in[x_{L},x_{R}]\to\left[\begin{array}[]{c}U^{*}(x)\\ H(x)\end{array}\right]$ (52)
define the same curve for all $x_{L}<x_{R}$. In particular, if the family of straight segments is chosen, then the numerical method is well-balanced for every stationary solution such that (52) lies in a straight line for every $x_{L}<x_{R}$. This property will be used in Section 5 to define numerical methods that preserve water-at-rest solutions for the two-layer shallow-water system. More sophisticated families of paths could be considered to preserve more general stationary solutions, as the ones based on the Generalized Hydrostatic Reconstruction introduced in [22] which will be done in a forthcoming paper.
The numerical treatment of the source term in this strategy can be interpreted as it was done in Section 2.2 for the nonconservative products $B(U)U_{x}$: the source term is first written as the derivative of a new flux function
$S(U)H_{x}=\partial_{x}\mathcal{S}^{L}_{i}$
with
$\mathcal{S}^{L}_{i}(x,t)=\int_{0}^{1}S(\Psi_{U}(s;W(x_{i},t),W(x,t))\partial_{s}\Psi_{H}(s;W(x_{i},t),W(x,t))\,ds,$
while in the global-flux approach, it is rewritten as
$S(U)H_{x}=\partial_{x}\mathcal{S}^{G}$
with
@@
Again, $\mathcal{S}^{L}_{i}$ is approximated at the node points of the stencil as follows
$\mathcal{S}^{L}_{i}(x_{j},t)\approx S_{\widetilde{\Psi}}(W_{i},W_{j})(H(x_{j})-H(x_{i})),\quad j\in\mathcal{S}_{i},$
and no integrals at the cells have to be approximated thus avoiding the need to calculate new reconstructions at the quadrature points.
3.2.2 Strategy 2
Strategy 2 extends to the nonconservative system the technique proposed in [44] for systems of balance laws. Unlike Strategy 1 the well-balanced property of the methods based on this strategy will not depend on either the choice of the family of paths or the matrix splitting.
Let us consider first the numerical method for (45) given by
$\frac{dU_{i}}{dt}+\frac{1}{\Delta x}\left(\widehat{D}^{-}_{i+1/2}+\widehat{D}^{+}_{i-1/2}\right)=S(U_{i})H_{x}(x_{i}),$ (53)
where the fluctuations $\widehat{D}^{+}_{i-1/2}$ are defined by (28) with any choice of family of paths, Roe matrix and matrix-splitting. Under the hypothesis of Proposition 2, this method is highly accurate but in principle does not preserve any stationary solution. Let us modify the method so that a given $m$-parameter family of stationary solutions
$U^{*}(x;c_{1},\dots,c_{m}),$ (54)
with $m\leq N$, is preserved. To do this, let us assume that there exists a $m\times N$ matrix $C$ with rank $m$ such that, given any point $\bar{x}$ and any state $\bar{U}$, there exists a unique stationary solution of the family satisfying
$CU^{*}(\bar{x})=C\bar{U},$ (55)
i.e., this system of equations determines the value of the $m$ parameters. The idea is then to rewrite the equation at $x_{i}$ equivalently as follows:
$\frac{dU_{i}}{dt}+A(U)U_{x}-A(U^{*}_{i})U^{*}_{i,x}=(S(U(x_{i},t))-S(U^{*}_{i}(x_{i},t)))H_{x}(x_{i}),$
where $U^{*}_{i}$ is the unique stationary solution of the $m$-parameter family satisfying
$CU_{i}^{*}(x_{i})=CU(x_{i},t).$ (56)
The idea is then to discretize $A(U)U_{x}$ and $A(U^{*}_{i})U^{*}_{i,x}$ together by applying the strategy introduced in Section 2.2 what leads to the numerical method:
$\frac{dU_{i}}{dt}+\frac{1}{\Delta x}\left(\widehat{D}^{-}_{i+1/2}+\widehat{D}^{+}_{i-1/2}\right)=(S(U_{i})-S(U^{*}_{i}(x_{i})))H_{x}(x_{i}),$ (57)
with
@@ 3.2.2 Strategy 2
$+\mathcal{R}^{R}(D^{-}_{i-k,i}-D^{*,-}_{i-k,i},\dots,D^{-}_{i+k,i}-D^{*,-}_{i+k,i}),$
where the starred fluctuations are given by
@@
The following result then holds.
@@
The numerical method (57)-(58)-(59) is well-balanced for all the stationary solutions of the family (54), i.e. $\{U^{*}(x_{i};c_{1},\dots,c_{m})\}$ is an equilibrium of the ODE system (19) for every stationary solution $U^{*}$.
The proof is straightforward. As an application, let us consider the family of stationary solutions
$U^{*}(x;c)=\left[\begin{array}[]{c}u^{*}(x;c)\\ v^{*}(x;c)\end{array}\right]=\left[\begin{array}[]{c}\frac{c}{2}+\sin(x)\\ \frac{c}{2}-\sin(x)\end{array}\right],\quad c\in\mathbb{R}$
of System (4). Given $\bar{x}$ and $\bar{U}=[\bar{u},\bar{v}]^{T}$ the equation
$u^{*}(x;c)+v^{*}(x;c)=\bar{u}+\bar{v}$
determines the value of the parameter
$c=\bar{u}+\bar{v},$
i.e. $m=1$ and $C=[1,1]$ in this case. Therefore, if the family of straight segments and the Roe matrix $A_{\Psi_{1}}$ in (27) are chosen, the numerical method (57) with
$U^{*}_{i}(x)=\left[\begin{array}[]{c}\frac{u_{i}+v_{i}}{2}+\sin(x)\\ \frac{u_{i}+v_{i}}{2}-\sin(x)\end{array}\right]$
is high-order accurate and well-balanced for the given family of stationary solution. This technique will be used in next section to design again a numerical method that preserves water-at-rest solutions for the two-layer shallow-water system that is based on the LF splitting.
For some particular problems, this strategy can be extended to design fully well-balanced methods, i.e. methods that preserve all the stationary solutions. In effect, let us suppose that, given $\bar{x}$ and $\bar{U}$, there is only one stationary solution $U^{*}(x)$ such that
$U^{*}(\bar{x})=\bar{U}$
(i.e. $m=N$ and $C$ is the identity matrix) or, there are several but it is possible to use a criterion to select one of them (like the flow regime, for instance). Then, the strategy can be applied by taking $U^{*}_{i}$ as the unique or the selected stationary solution such that
$U^{*}_{i}(x_{i})=U(x_{i},t).$
In this case, the numerical method reduces to (19), with the fluctuations given by (58)-(59).
In particular, this technique allows the design of fully well-balanced numerical methods for the shallow-water system: see [44]. In fact, the numerical methods introduced here based on Strategy 2 reduce to the ones in this reference when they are applied to systems of balance laws, as the shallow-water system. In the reference, it has been shown that theses methods deal correctly with discontinuous bottom functions $H$: the fully-well balanced methods are able to correctly capture the stationary contact discontinuities standing on the points of discontinuity of $H$: see [44] for details. This technique can be extended to derive numerical methods that are fully well-balanced for the 1D two-layer shallow-water systems but the computation of moving stationary solution is more involved: this will be the object of future work.
3.3 Implementation
We summarize here the implementation of two well-balanced methods: Method 1 in which Strategy 1 is combined with the Upwind splitting, and Method 2 corresponding to Strategy 2, LF splitting, and reconstruction in characteristic variables. Let us assume that approximations $U_{j}$ to the solution are available at the cell points; the fluctuations and the source terms at a point $x_{i}$ are then computed as follows;
  1. 1.
    Compute the Roe matrices $A_{i,j}=A_{\Psi}(U_{i},U_{j})$, $j=i-k,\dots,i+k$ and their eigenvalues $\{\lambda_{i,j;l}\}_{l=1}^{N}$.
  2. 2.
    Compute the fluctuations at the stencil points as follows:
    1. (a)
      Method 1:
      Compute the eigenvector matrices $R_{i,j}$, $j=i-k,\dots,i+k$
      Solve the linear systems
      $R_{i,j}\vec{\alpha}=U_{j}-U_{i}-A_{i,j}^{-1}S_{i,j}(H(x_{j})-H(x_{i}))$
      to obtain $\{\alpha_{i,j;l}\}_{l=1}^{N}$ such that
      $U_{j}-U_{i}-A_{i,j}^{-1}S_{i,j}(H(x_{j})-H(x_{i}))=\sum_{l=1}^{N}\alpha_{i,j;l}\vec{r}_{i,j;l},$
      Compute then
      $D_{i,j}^{\pm}=\sum_{i=1}^{N}\alpha_{i,j;l}\lambda^{\pm}_{i,j;l}\vec{r}_{i,j;l}.$
    2. (b)
      Method 2:
      Compute first the stationary solution $U^{*}_{i}$ of the family (54) satisfying
      $CU^{*}_{i}(x_{i})=CU_{i}$
      and evaluate it at $x_{j}$, $j=i-k,\dots,i+k$.
      Compute the Roe matrices $A^{*}_{i,j}=A_{\Psi}(U_{i}^{*}(x_{i}),U^{*}_{i}(x_{j}))$.
      Compute then
      $D^{\pm}_{i,j}$ $=\dfrac{1}{2}\left(A_{i,j}\left(U_{j}-U_{i}\right)\pm\alpha\left(U_{j}-U_{i}\right)\right),$
      $D^{*,\pm}_{i,j}$ $=\dfrac{1}{2}\left(A^{*}_{i,j}\left(U_{i}^{*}(x_{j})-U_{i}^{*}(x_{i})\right)\pm\alpha\left(U_{i}^{*}(x_{j})-U_{i}^{*}(x_{i})\right)\right),$
      where $\alpha$ is chosen so that $\alpha>|\lambda_{i,j;l}|$ for all $i,j,l$.
  3. 3.
    Compute the WENO reconstructions of the fluctuations at the cell interfaces as follows:
    1. (a)
      Method 1:
      $\widehat{D}^{-}_{i+1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i,i-k},\dots,D^{+}_{i,i+k})+\mathcal{R}^{R}(D^{-}_{i,i+1-k},\dots,D^{-}_{i,i+1+k}),$
      $\widehat{D}^{+}_{i-1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i-1-k,i},\dots,D^{+}_{i-1+k,i})+\mathcal{R}^{R}(D^{-}_{i-k,i},\dots,D^{-}_{i+k,i}).$
    2. (b)
      Method 2:
      Compute the matrices of right and left eigenvectors, $R_{i-1,i}$, $L_{i-1,i}$, (resp. $R_{i,i+1}$, $L_{i,i+1}$) of $A_{i-1,i}$ (resp. $A_{i,i+1}$)
      Then compute:
      $\widehat{D}^{-}_{i+1/2}$ $=$ $R_{i,i+1}\mathcal{R}^{L}(L_{i,i+1}\left(D^{+}_{i,i-k}-D^{*,+}_{i,i-k}\right),\dots,L_{i,i+1}\left(D^{+}_{i,i+k}-D^{*,+}_{i,i+k}\right))$
      $+R_{i,i+1}\mathcal{R}^{R}(L_{i,i+1}\left(D^{-}_{i,i+1-k}-D^{*,-}_{i,i+1-k}\right),\dots,L_{i,i+1}\left(D^{-}_{i,i+1+k}-D^{*,-}_{i,i+1+k}\right)),$
      $\widehat{D}^{+}_{i-1/2}$ $=$ $R_{i-1,i}\mathcal{R}^{L}(L_{i-1,i}\left(D^{+}_{i-1-k,i}-D^{*,+}_{i-1-k,i}\right),\dots,L_{i-1,i}\left(D^{+}_{i-1+k,i}-D^{*,+}_{i-1+k,i}\right))$
      $+R_{i-1,i}\mathcal{R}^{R}(L_{i-1,i}\left(D^{-}_{i-k,i}-D^{*,-}_{i-k,i}\right),\dots,L_{i-1,i}\left(D^{-}_{i+k,i}-D^{*,-}_{i+k,i}\right)).$
      Compute finally the source term:
      $\left(S(U_{i})-S(U^{*}_{i}(x_{i}))\right)H_{x}(x_{i}).$
Remark 2.
Although in Section 2.2 it was said that the numerical methods based on the Upwind splitting are more computationally expensive than those based on the LF splitting, observe that the numerical treatment of the source term in Method 2 requires the computation of the fluctuations corresponding to the chosen stationary solutions so that, as it will be seen in Section 6, the computational costs of both methods are comparable.